OpenStudy (anonymous):

Please if I have the following equation y=e^x, and before taking log I need to add 1 to both sides so the eq is y+1=e^x + 1 how do I apply logarithms and linearize the equation?

OpenStudy (anonymous):

applying Ln\[\ln y =\ln e^x\]\[\ln y =x\]

OpenStudy (anonymous):

y=e^x log both sides, log y = log e^x the x you can bring it over to the front hence, log y = x log e so compare to y=mx + c you can think y similar to log y, x is still x, log e is the gradient of the line and c=0 I think there is no use of adding 1 on both sides as it's getting you nowhere.

OpenStudy (anonymous):

well, I must add 1 before to transform my variable into a normal distribution that's the reason of adding one.. but anyway I already solved! !:) thanks a lot!!